# Policies for multiplexed quantum repeaters: theory and practical performance analysis

### v1

@@ III Policies for multiplexing
When multiple links are available between nodes of the network, there are many possibilities for how to perform entanglement swapping. There is also the possibility to perform entanglement distillation between nodes that share several links. In linear quantum networks without multiplexing (i.e., $n_{ch}=1$ in Fig. 1), one of the simplest and best studied swapping protocols is swap-as-soon-as-possible (swap-asap). As the name suggests, the swap-asap policy dictates that entanglement swapping should be attempted as soon as two memories in a given node become active, i.e., as soon as an entanglement swapping policy becomes possible. swap-asap policies have been shown to be better than fixed nesting or doubling policies for a large set of parameter regimes in linear chain networks consisting of a single channel between nodes [30, 25, 75]. The policies we present in this work are generalizations of the swap-asap policy to the multiplexing setting.
The direct application of swap-asap in multiplexed linear chains is unspecified due to the increased degrees of freedom. In particular, two immediate questions arise:
@@
    If the opportunity is available simultaneously, should we preferably perform entanglement swapping on links that result in a virtual link between farthest nodes or the ones that result in a strongest—highest fidelity—virtual link? Based on this question, we define the following two generalizations of swap-asap for multiplexed linear chains.
@@
      The Farthest Neighbor (FN) swap-asap policy: This policy prioritizes the creation of virtual links between faraway nodes. Accordingly, we rank the links based on their lengths (farthest links are ranked first), and then entanglement swapping is performed in that order. This policy is illustrated with an example in Fig. 2.
@@
      The Strongest Neighbor (SN) swap-asap policy: This policy prioritizes the creation of virtual links with high fidelity. Accordingly, we rank the links based on their ages (links with the lowest age are ranked first), and then entanglement swapping is performed in that order.
@@
Throughout the rest of this work, we demonstrate via simulations that our quasi-local SN and FN swap-asap multiplexing policies can outperform (in terms of waiting time and end-to-end fidelity) not only non-multiplexed or parallel policies but also fully-local multiplexed policies in which decisions about entanglement swapping are made without utilising any knowledge of the network’s state,
like in the random multiplexing policy. We also demonstrate via simulations that our policies can outperform the well-known and widely-used doubling and nested-purification-and-swapping protocol.
@@
IV Performance of multiplexing based policies
Having defined our SN and FN swap-asap policies in the previous section, we now evaluate their performance. We consider two figures of merit: the average waiting time and the average age of the youngest end-to-end link. In this section, we measure the average waiting time in terms of the number of time steps. This unitless, abstract quantity can be converted into physical times, and also the ages converted to fidelity values, by incorporating the network hardware and design parameters. Examples of such a translation are given in Sec. IV and further details are provided in Sec. VII.1 and Appendix A.
In Fig. 3, and for all of the plots shown in this paper, the parameter values that we choose are based on what is computationally feasible, but they are also based on what we expect with actual hardware, as we outline in Sec. VII.1. The choice of network parameters such as $p_{\ell},m^{\star}$ etc. in specific examples is also motivated by the intent to show the most prominent and overarching trends.
Of course, FN, SN, and random are not the only three possible options for multiplexing-based policies. For example, one could develop a ranking scheme based on a combination of age and distance, or one could perform a search over all possible pairings of links, looking for the optimal pairings in terms of some figure of merit (such as the sum of lengths of all virtual links formed). With the respect to such optimal ranking schemes, it is not clear how close the SN and FN schemes are to being optimal; they are, however, simple and intuitive policy choices that require no optimization in terms of pairing of links for entanglement swapping, and hence are practical for near-term implementation. We defer the discussion of other policy alternatives and comments on the optimality of SN and FN swap-asap to Appendix D.
IV.1 Comparison to doubling
@@
Let us now compare our policies with those in existing literature. In particular, we compare our policies with one widely studied in previous works, namely, the doubling (or nesting) policy; see, e.g., Refs. [15, 16, 58, 59, 64, 79]. In this policy, which only applies to repeater chains with $2^{N}$ elementary links, the lengths of links are always doubled. As an example, consider the following situation. Suppose two adjacent elementary links are formed, an entanglement swap is performed successfully, and a link of length two is obtained. Now, in contrast to the swap-asap policy, this length-two link cannot be extended by swapping it with links of arbitrary lengths—links of length two must be swapped with links of length two only—which means that we must wait for another link of length two to be created before an entanglement swap can be attempted. In the meantime, the first produced virtual link keeps aging. We can thus see that this doubling policy is more restricted compared to swap-asap. Accordingly, it was shown in Ref. [75] that for low elementary link success probabilities and moderate swapping success probabilities (such as 50%, as in the case of linear optics), the swap-asap policy outperforms the doubling policy, although infinite coherence times for memories was assumed in that work.
@@
Figure 7: Average waiting time (top) and average age of the youngest end-to-end link (bottom) for a repeater chain with $n=9$ nodes (eight elementary links), according to the FN, SN, and random swap-asap policies, compared to the SN doubling policy. For small values of $p_{\ell}$, both the FN and SN swap-asap policies outperform the doubling policy in terms of waiting time. The SN swap-asap policy outperforms doubling for the average age for all values of $p_{\ell}$ up to $p_{\ell}=0.5$.
Now, within the multiplexing scenario, the doubling policy can be implemented by ranking links based on age when more than one option of entanglement swapping is available. Thus, we call this policy Strongest-Neighbor doubling or SN doubling. Our aim now is to see if, within the multiplexing setting, and with finite coherence times for the quantum memories, our FN and SN swap-asap policies can outperform a doubling based multiplexing policy. This is what we show in Fig. 7. In terms of the average waiting time, the FN swap-asap policy outperforms the SN doubling policy for all values of link success probability $p_{\ell}$ up to $p_{\ell}=0.5$, including very low values of $p_{\ell}$, which is often most relevant in practice. It is important to mention here that this result is not true for all parameter regimes; for high values of $p_{\ell}$ and $m^{\star}$, the doubling policy can actually outperform our swap-asap policies. Our results are consistent with the results in previous works. Specifically, our results generalize those in Ref. [75] to the case of finite coherence times.
At the same time, our results also show the importance of ranking the links in an appropriate way, reflected in the fact that the random swap-asap policy and the SN swap-asap policies actually perform worse than the doubling policy in terms of average waiting time. We see very similar results in terms of the average age of the youngest end-to-end link. With respect to this figure of merit, the SN swap-asap policy outperforms all other policies discussed here, including the SN doubling policy. Below, in Sec. V, we also include distillation in our swap-asap policies and the doubling policy to compare their performance, since the latter has also been widely discussed in existing literature; see, e.g., Refs. [15, 16, 58, 59, 64, 79]. The doubling policy is also quasi-local in the same sense as SN and FN swap-asap, because a node needs to know the length of its links, i.e., it needs to know which nodes it is connected to. This is an important consideration when accounting for classical communication overheads of this policy, which we do in Sec. VI.
V Multiplexing policies with entanglement distillation
We now consider entanglement distillation with the probabilistic BBPSSW protocol [53], which allows for distillation of two entangled links to one. We provide a summary of the protocol in Appendix B. Briefly, if $f_{1}$ and $f_{2}$ are the fidelities of the two links being distilled, then the success probability $P_{\text{distill}}(f_{1},f_{2})$ of the protocol and the resulting fidelity $F_{\text{distill}}(f_{1},f_{2})$ are given by
The purpose of this section is to assess the usefulness of entanglement distillation for quantum repeater chains with multiple memories per repeater. In particular, in this section, we examine the extent to which entanglement distillation can compensate for the loss of fidelity incurred due to entanglement swapping and waiting (decoherence).
V.1 Initial considerations
Recall that entanglement swapping of two links of ages $m_{1}$ and $m_{2}$ results in a virtual link of age $m_{1}+m_{2}$, and waiting increases the ages of all active links by one time unit. Similarly, as we show in Appendix B, the age after distillation with the BBPSSW protocol is given by the following formula:
$m^{\prime}=\left\lceil m^{\star}\log\left(\frac{15-6(f(m_{1})+f(m_{2}))+24f(m_{1})f(m_{2})}{32f(m_{1})f(m_{2})-2(f(m_{1})+f(m_{2}))-1}\right)\right\rceil,$ (4)
where $m_{1}$ and $m_{2}$ are the ages of the two links being distilled, the function $f(m)$ is defined in Eq. (1), and $\lceil x\rceil$ is the smallest integer greater than or equal to $x$.
@@ V.1 Initial considerations
Another important consideration is that both links being distilled should be entangled. If one of the links is not entangled, then it is better to simply discard the unentangled link and keep the entangled one. We elaborate on this point in Appendix B. With this in mind, let us note that under the Pauli noise model that we consider, the fidelity according to Eq. (1) is bounded from below by $0.25$ (in the limit $m\to\infty$), and it is equal to approximately $0.5259$ when $m=m^{\star}$. The entanglement threshold for the states that we consider is $0.5$, meaning that the link is entangled if and only if its fidelity strictly exceeds $0.5$, and thus the age at which the link is no longer entangled is $m=\lceil m^{\star}\log(3)\rceil\approx\lceil 1.09m^{\star}\rceil$; see Appendix A for the details. This means that, depending on the value of $m^{\star}$, it can be useful to distill links that are older than $m^{\star}$ time steps. However, the probability of successful distillation falls with increasing ages of the links, and hence using a very old link in conjunction with a very young link increases the likelihood of wasting the high-quality young link. Thus, while distillation might be useful in this scenario, it is not necessarily worthwhile. Also, for the moderate values of $m^{\star}$ considered in our simulations, we choose to discard links at $m^{\star}$ time steps, instead of the strict limit of $\lceil m^{\star}\log(3)\rceil$, even when distillation is considered.
V.2 Distillation policies
  • Pumping policy: Referred to as “scheme C” in Ref. [16], this policy proceeds as follows. Two links of fidelity $f_{0}$ are distilled, then a fresh link of fidelity $f_{0}$ is used to distill the previously distilled link, and so on. This policy is tailored to a practical situation in which fresh links are used to distill the surviving links of previous time steps, and thus the links being distilled do not necessarily have the same fidelity, as in the banded policy. If the initial fidelity satisfies $f_{0}>\frac{1}{2}$, the fidelity $f_{r}$ after $r\in\{1,2,\dotsc\}$ rounds of the pumping policy is given by the recurrence relation $f_{r}=F_{\text{distill}}(f_{r-1},f_{0})$. In Appendix B.1, we provide a closed-form, analytical solution to this recurrence relation, and using that solution we show that in the limit of arbitrarily many rounds of pumping, the fidelity saturates as follows:
    $\lim_{r\to\infty}f_{r}=\frac{6f_{0}-3+\sqrt{7-26f_{0}+28f_{0}^{2}}}{8f_{0}-2}.$ (5)
    The limiting fidelity on the right-hand side of this equation is strictly less than one whenever $f_{0}<1$. This proves that, in contrast to the banded policy, with the pumping policy it is not possible to distill links to arbitrarily high fidelity, even with an infinite number of rounds.
  • Distill as soon as possible: This is a sequential policy, which we introduce in this work as a hybrid of the banded and pumping policies, and very similar in spirit to the greedy policy defined in Ref. [79]. We sort the links in increasing order of their ages, and then pair them in increasing order. After one round of pairwise distillation, successfully distilled links are again sorted age-wise and the distillation is attempted again, until only one distilled link or no active link remains. This approach therefore uses the benefits of the banded policy, by pairing links of very similar ages, and at the same time, by realising the restrictions introduced by decoherence and finite number of channels, by pumping fresh links into the queue of available links as soon as possible. We therefore refer to this policy as “distill-as-soon-as-possible” (distill-asap), and this is the policy that we consider for our simulations in Sec. V.3 below.
Now, how should we combine entanglement distillation with entanglement swapping? Should we distill first then swap—a policy we call distill-swap—or the other way around, which we refer to as swap-distill? In Appendix B.2 and Appendix C, we provide analytical calculations that compare the two policies for a three-node chain. For the rest of this section, we compare the policies with full-fledged simulations of long quantum repeater chains.
V.3 Distill-swap vs. swap-distill: simulation results
Having developed some intuition about the performance of the distillation based policies through analytical results, let us now consider the fully dynamic setting with long quantum repeater chains. Here, the policy that we use is the distill-asap policy defined in Sec. V.2 above. We stress again that the aim here is not to show optimality of our approach. The pairing of links for distillation is an optimization problem in its own right and not the focus of this work.
@@ V.2 Distillation policies
We now present simulation results for a five-node chain with five channels. A somewhat large value of the cutoff is chosen ($m^{\star}=24$) to show some important scaling behaviour with increasing age $m_{0}$ of the fresh (newly created) links. Fig. 8 shows the average waiting time and average age of end-to-end links as a function of the ages of the elementary links $m_{0}$ when they are fresh, i.e., newly created. For both swap-distill and distill-swap, we have fixed the entanglement swapping policy to FN swap-asap.
@@
It is clear that swap-distill is the policy of choice in terms of reducing the waiting time when fresh elementary links have high fidelity (low age); otherwise, distill-swap performs better. At the same time, distill-swap always performs better in terms of the age of the end-to-end link. This result reinforces our intuition, developed from the previous analytical results. We also verify that waiting time rapidly increases beyond a threshold value of $m_{0}$ for both swap-distill and distill-swap, the threshold being smaller for swap-distill. Beyond the $m_{0}$ values reported in Fig. 8, end-to-end entanglement generation time increases so much in the simulations that we are unable to find any reasonable estimate of an average value. Thus, for all practical purposes, the waiting time is essentially infinite. This is also indicated by the average ages approaching $m^{\star}$.
Some estimates of the thresholds up to which the two distillation policies are effective can be made using some rather simple arguments. If the fresh elementary links have age $m_{0}$, the age of the distilled link saturates to a minimum value $m^{d}_{min}(m_{0})$ which is a function of $m_{0}$. In order to get an end-to-end link via entanglement swapping of such distilled links, the following condition must be satisfied:
$(n-1)\cdot m^{d}_{min}(m_{0})\leq m^{\star}.$ (6)
For the configuration in Fig. 8, the above condition translates to $m^{d}_{min}\lessapprox 4$, which in terms of $m_{0}$ is $m_{0}\leq 5$, and this is the value around which we see that the average waiting time diverges.
In the case of swap-distill, a typical trajectory to get an end-to-end link would be to first perform $(n-2)$ swaps to get links of age $(n-1)\cdot m_{0}$, which are then distilled to links of age $m^{d}_{min}((n-1)\cdot m_{0})$. The threshold condition thus takes the following form:
$m^{d}_{min}((n-1)\cdot m_{0}))\leq m^{\star},$ (7)
but since distillation is ineffective when both links involved are above the cutoff age, this condition is equivalent to
@@
This also explains why the threshold is lower for swap-distill, since $m^{d}_{min}(m_{0})\leq m_{0}$. Again, for the configuration in Fig. 8, this translates to $m_{0}\lessapprox 4$, which agrees well with the simulation results. Furthermore, the threshold in the case of the swap-distill policy is identical to a policy without distillation.
@@
We further elaborate on our answer to the question of whether to distill first or to swap first in Fig. 9. We plot the waiting time improvement factor for distillation, which we define as the ratio of the average waiting time without distillation to the average waiting time with distillation. Apart from the quality of fresh links, as shown in Fig, 8, the choice between the two distillation policies also depends on the other network parameters. As an illustration, we show that for the parameters ($n,n_{ch},m^{\star},p_{\ell}$) chosen in Fig. 9, when $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. This can be understood in light of analytical results in Appendix C and Fig. 8. In the regime of low $p_{\ell}$, links are on average older and we have seen in the previous discussions in this section that distill-swap performs better in this case since long entanglement swaps can only be possible once the ages of the participating links have been reasonably reduced; also, distilling after entanglement swapping is highly improbable to be successful because the ages are going to only increase due to the swapping. On the other hand, when $p_{\ell}$ is high, swap-distill performs better, because links are on average younger and much more conducive to successful entanglement swapping. Further we see that for large $p_{\ell}$ distillation is either not very useful or even detrimental. Therefore, the decision to distill or not to distill also should be taken after considering the network parameters. We explore this point further in Fig. 10.

### v2

@@ III Quasi-local multiplexing policies
We now introduce our two multiplexing policies for entanglement distribution in a linear chain quantum network. In linear quantum networks without multiplexing (i.e., $n_{ch}=1$ in Fig. 1), one of the simplest and best studied swapping protocols is swap-as-soon-as-possible (swap-asap). As the name suggests, the swap-asap policy dictates that entanglement swapping should be attempted as soon as two memories in a given node become active, i.e., as soon as an entanglement swapping policy becomes possible. The swap-asap policy has been shown to be better than fixed nesting or doubling policies for a large set of parameter regimes in linear networks consisting of a single channel between nodes [31, 26, 56].
The direct application of swap-asap in multiplexed linear chains is unspecified due to the increased number of degrees of freedom. Indeed, when multiple links are available between nodes of the network, there are many possibilities for how to perform entanglement swapping. There is also the possibility to perform entanglement distillation between nodes that share several links. We therefore define our policies based on the following considerations.
@@
    If entanglement swapping of multiple pairs of links is possible, how should we group the pairs? We could, for example, perform entanglement swapping on link pairs that result in virtual links between farthest nodes, or the ones that result in strongest—highest fidelity—virtual links. Based on this consideration, we define the following two generalizations of swap-asap for multiplexed linear chains.
@@
      The Farthest Neighbor (FN) swap-asap policy: This policy prioritizes the creation of virtual links between faraway nodes. Accordingly, we rank the links based on their lengths (farthest links are ranked first), and then entanglement swapping is performed by pairing the links with the same rank. This policy is illustrated with an example in Fig. 2. Intuitively, this policy prioritizes minimizing the average waiting time.
@@
      The Strongest Neighbor (SN) swap-asap policy: This policy prioritizes the creation of virtual links with high fidelity. Accordingly, we rank the links based on their ages (links with the lowest age are ranked first), and then entanglement swapping is performed in that order. Intuitively, this policy aims to create the strongest or highest-fidelity links between the end nodes of the network.
@@
    How should we distill multiple links to one? For this, we introduce the distill-as-soon-as-possible (distill-asap) policy. This is a sequential policy, which can be thought of as a hybrid of the banded and pumping policies [17] (see also Appendix C), and very similar in spirit to the greedy policy defined in Ref. [84]. We sort the links in increasing order of their ages, and then pair them in increasing order. After one round of pairwise distillation, successfully distilled links are paired again and distillation is attempted again, until only one distilled link or no active link remains. This approach therefore uses the benefits of the banded policy, by pairing links of very similar ages, and at the same time, by realising the restrictions introduced by decoherence and finite number of channels, by pumping fresh links into the queue of available links as soon as possible.
  3. 3.
@@
The FN and SN swap-asap policies that we have introduced have classical communication (CC) overheads. Indeed, whenever an entanglement swapping decision is made, all the memories at the node are assumed to have knowledge of which nodes/memories they are connected to. This information is needed to decide the order of entanglement swaps based on age (SN) or length (FN) of the links. Thus, every node has some “quasi-local” knowledge of the network state but not the “global” network state.
We demonstrate via simulations that our quasi-local SN and FN swap-asap multiplexing policies can outperform (in terms of waiting time and end-to-end fidelity) the non-multiplexed, parallel swap-asap policy (as we would expect), and also the random swap-asap policy. We present these results in Appendix A. In the next section, we present results that demonstrate, more importantly, that our policies can outperform fully local multiplexing policies, as well as the well-known and widely-used nested-purification-and-doubling protocol.
@@
IV Performance evaluation
We now evaluate the performance of our policies from Sec. III with respect to the average waiting time and average age of the end-to-end link.
Figure 3: Average waiting time (top) and average age of an end-to-end link (bottom) as a function of the elementary link success probability when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy. The FN swap-asap policy provides smaller values for both figures of merit. Similar trends can also be seen for the SN swap-asap policy.
In order to benchmark our quasi-local policies fairly against fully-local policies, some modifications need to be made to the way the network evolves under the fully-local policies. These modifications are necessary, because when entanglement swaps are probabilistic, if links need to be restarted after a failed Bell state measurement, but before their actual memory cutoff, some CC will be needed. Therefore, if a policy has to be fully-local and free of any CC overheads, nodes must be agnostic to entanglement swapping failures. All nodes, once they have a heralded entangled link, must retain them until their anticipated memory cutoff age $m^{\star}$. Once this restriction is made, the only available information to nodes that can be used to decide the rank of states is the perceived or local ages of its memories. In other words, every link in the network now has two ages: one is its real age, as determined by probabilistic swaps in the evolution “in real-time”; the second is its perceived age, which is the age that the nodes that hold the link assume without any knowledge of the outcomes of entanglement swaps. When the perceived ages are used to make decisions in the network, there is no CC overhead during the evolution. We call such a policy local age-based (ranking) multiplexed swap-asap policy. For such a policy, the CC cost is given by only one round of end-to-end communication time, in the last time step.
From Fig. 3, it is clear that substantial advantage for both the waiting time (around two orders of magnitude reduction) and fidelity is obtained by using our policies compared to the local age-based swap-asap policy, even when CC times are taken into consideration.
Figure 4: (Top) Average waiting time as a function of the number of nodes when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy, and we find that the FN swap-asap policy provides smaller waiting times. More surprisingly, the advantage grows with an increasing number of nodes. This provides evidence of the scalability of the advantage gained by using knowledge of the network state. (Bottom) Number distribution of lengths of entanglement swaps in the FN swap-asap policy. We observe an exponential suppression of entanglement swaps of the order of the length of the chain, which is an important reason for the retained advantage of this policy, even with classical communication included.
A more surprising observation is the fact that even as the number of nodes increases, the advantage of FN swap-asap over the local age-based ranking policy remains strong—in fact, the advantage increases with increasing number of nodes. In Fig. 4 (top), we show the average waiting times for the local age-based ranking and the FN swap-asap policies as a function of the number of nodes. Intuitively, we might have guessed that as the number of nodes increases, so too would the CC cost, because longer and longer swaps would be required to distribute entanglement between the end points of the chain, making quasi-local policies worse off in terms of average waiting times. It is therefore surprising that not only a large improvement using the FN and SN swap-asap policies is obtained for long chains, but further that the improvement increases with the growing scale (number of nodes) of the network.
Nonetheless, we can explain this non-intuitive result as follows. First, we remind the reader that although fully-local policies have the advantage of no CC overheads, they suffer in terms of average waiting time due to the agnostic entanglement swapping. Indeed, waiting to restart links that are already dead is detrimental to the average waiting time. Second, because we use quasi-local policies, the CC costs in a given time step are upper-bounded by the length of the longest entanglement swapping operation, rather than the length of the entire repeater chain. (Recall that we define the length of an entanglement swapping operation as the length, in terms of number of nodes, of the longer of two links involved in the swap.) In a typical time-evolution of a network, from completely disconnected to an end-to-end connected link, for the FN swap-asap policy the number of entanglement swapping operations of length equal to the number of nodes is roughly $O(1)$, i.e., roughly constant with respect to the number of nodes. This is seen in Fig. 4 (bottom). The average number distribution of the length of entanglement swaps shows an exponentially suppressed tail. Therefore, there are very few time steps in which the CC cost is given by the entire length of the repeater chain. This justifies and supports the use of quasi-local policies over global policies for large networks, because in the latter, the CC cost is equal for each time step and given by the entire length of the repeater chain, while for quasi-local policies the CC cost can be considerably less. Also see Appendix F for details.
IV.1 Comparison with doubling
@@
Let us now compare our policies with those in existing literature. In particular, we compare our policies with one widely studied in previous works, namely, the doubling (or nesting) policy; see, e.g., Refs. [16, 17, 61, 62, 67, 84]. In this policy, which only applies to repeater chains with $2^{N}$ elementary links, the lengths of links are always doubled. As an example, consider the following situation. Suppose two adjacent elementary links are formed, an entanglement swap is performed successfully, and a link of length two is obtained. Now, in contrast to the swap-asap policy, this length-two link cannot be extended by swapping it with links of arbitrary lengths—links of length two must be swapped with links of length two only—which means that we must wait for another link of length two adjacent to the original link to be created before an entanglement swap can be attempted. In the meantime, the first produced virtual link keeps aging. We can thus see that this doubling policy is more restricted compared to swap-asap. Accordingly, it was shown in Ref. [56] that for low elementary link success probabilities and moderate swapping success probabilities (such as 50%, as in the case of linear optics), the swap-asap policy outperforms the doubling policy, although infinite coherence times for memories was assumed in that work.
@@
Figure 5: Comparison of (top) average waiting times and (bottom) average age of the end-to-end link for the FN and SN swap-asap policies and the SN doubling policy. FN swap-asap outperforms both SN swap-asap and SN doubling in terms of average waiting time, and SN swap-asap performs the best in terms of average age of the end-to-end link, for all elementary link success probabilities ($p_{\ell}$).
Now, within the multiplexing scenario, the doubling policy can be implemented by ranking links based on age when more than one option of entanglement swapping is available. Thus, we call this policy Strongest-Neighbor doubling, or SN doubling. This doubling policy is also quasi-local, in the same sense as SN and FN swap-asap, because a node only needs to know the length of its links, i.e., it needs to know which nodes it is connected to.
Our aim now is to see if, within the multiplexing setting, and with finite coherence times for the quantum memories, our FN and SN swap-asap policies can outperform the SN doubling policy. Our results are shown in Fig. 5. We see that FN swap-asap outperforms both SN swap-asap and SN doubling in terms of average waiting time. At the same time, SN swap-asap performs better than SN doubling, thus further strengthening our intuition that swap-asap policies outperform doubling policies in practical resource-constrained scenarios. We also see that SN swap-asap performs the best in terms of average age of the end-to-end link, for all elementary link success probabilities $p_{\ell}$. This is expected, because SN swap-asap prioritizes the creation of younger links.
Figure 6: Comparison of (top) average waiting times and (bottom) average age of the end-to-end link, for the FN swap-asap policy with and and without distillation. For the average waiting time, distill-swap outperforms the policy without distillation for low elementary link success probabilities, but it is not useful to distill when the success probability is high. At the same time, we see that it is never useful to swap-distill, both in terms of the average waiting time and the age of the end-to-end link. The CC costs of distillation are much higher when long virtual links are distilled, which provides intuition for the inefficacy of the swap-distill policy.
IV.2 Distillation-based policies
Now we present results for distillation-based policies when CC overheads are included. As mentioned earlier in Sec. II.1, the CC cost of distillation for any quasi-local policy is given by the length of longest links that are used for distillation. For local policies, distillation is only allowed at the elementary link level, and only when the links are freshly prepared. We continue using the distill-asap policy described in Sec. III. The CC overheads make the already probabilistic BBPSSW protocol even more costly. Furthermore, the CC costs of distilling longer (virtual) links is higher than distilling elementary links, and thus the question of whether to distill or not and whether to swap first or to distill first depends crucially on the CC overheads. We show the average waiting times for FN swap-asap policy without distillation and both distillation-based policies (swap-distill and distill-swap) in Fig. 6.
Figure 7: Comparison of (top) average waiting times and (bottom) average age of end-to-end link, distill-swap versions of FN swap-asap, SN swap-asap and the SN doubling policies when classical communication costs are included. FN swap-asap is the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
Finally, in Fig. 7, we again compare the different distillation-based versions of quasi-local policies amongst themselves, viz., FN swap-asap, SN swap-asap and the SN doubling policies. We see that, with distillation (distill-swap version) and the relevant CC costs added, the trends remain similar to Fig. 5, i.e., FN swap-asap is still the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
The results of the numerical simulations presented in this section show that including entanglement distillation in multiplexing policies introduces many new considerations, such as the fact that, in some parameter regimes, it is better to distill first and in others the other way around. In the next section, we take a deeper dive into these factors surrounding the use of entanglement distillation.
V Unraveling the role of entanglement distillation policies for multiplexing
The BBPSSW entanglement distillation protocol considered in this work non-deterministically converts two links with low fidelity into one link with a higher fidelity. This immediately raises the following question: “Is distillation useful for all network parameter regimes?” Due to its non-deterministic nature, in certain parameter regimes, the slight increase in fidelity obtained by distillation might not be worth the potential loss of links, in case it fails. Furthermore, even without the consideration of CC costs, there exists an underlying asymmetry between the way entanglement swapping and distillation protocols work. In the former, the success probability is independent of the ages of the links, and thus it does not discriminate between longer (generally older) and shorter links, whereas the latter’s success crucially depends on the ages of the input links. This asymmetry leads to the the following question: “Should we distill and then swap or the other way around?” This consideration leads to two distinct policies, namely distill-swap and swap-distill, with different behaviours in different parameter regimes, as we show in Fig. 6. The aim of this section is to pursue the mechanisms behind these different behaviours and unravel them layer by layer. We make one simplifying assumption in this section, which is that classical communication costs associated with entanglement swapping and distillation are ignored. This is motivated by two reasons: firstly, for short chains, CC costs are usually ignored in existing literature on entanglement distribution policies [44, 86, 87, 88, 31], thus this assumption allows us to fairly compare the performance of our policies with those in the existing literature. Secondly, ignoring CC simplifies our model and makes some analyses more transparent. At the same time, the conclusions we come to in this section hold qualitatively even when CC is accounted for.
@@ V Unraveling the role of entanglement distillation policies for multiplexing
Another important consideration is that both links being distilled should be entangled. If one of the links is not entangled, then it is better to simply discard the unentangled link and keep the entangled one. We elaborate on this point in Appendix C. With this in mind, let us note that under the Pauli noise model that we consider, the fidelity according to Eq. (1) is bounded from below by $0.25$ (in the limit $m\to\infty$), and it is equal to approximately $0.5259$ when $m=m^{\star}$. The entanglement threshold for the states that we consider is $0.5$, meaning that the link is entangled if and only if its fidelity strictly exceeds $0.5$, and thus the age at which the link is no longer entangled is $m=\lceil m^{\star}\log(3)\rceil\approx\lceil 1.09m^{\star}\rceil$; see Appendix B for the details. This means that, depending on the value of $m^{\star}$, it can be useful to distill links that are older than $m^{\star}$ time steps. However, the probability of successful distillation falls with increasing ages of the links, and hence using a very old link in conjunction with a very young link increases the likelihood of wasting the high-quality young link. Thus, while distillation might be useful in this scenario, it is not necessarily worthwhile. Also, for the moderate values of $m^{\star}$ considered in our simulations, we choose to discard links at $m^{\star}$ time steps, instead of the strict limit of $\lceil m^{\star}\log(3)\rceil$, even when distillation is considered.
@@
We now present simulation results for a seven-node chain with seven channels. (We also provide analytical results in Appendix D.) A somewhat large value of the cutoff is chosen ($m^{\star}=24$) to show some important scaling behaviour with increasing age $m_{0}$ of the fresh (newly created) links. Fig. 8 shows the average waiting time and average age of end-to-end links as a function of the ages of the elementary links $m_{0}$ when they are fresh, i.e., newly created. For both swap-distill and distill-swap, we have fixed the entanglement swapping policy to FN swap-asap.
@@
It is clear that swap-distill is the policy of choice in terms of reducing the waiting time when fresh elementary links have high fidelity (low age); otherwise, distill-swap performs better. At the same time, distill-swap always performs better in terms of the age of the end-to-end link. We also verify that the waiting time rapidly increases beyond a threshold value of $m_{0}$ for both swap-distill and distill-swap, the threshold being smaller for swap-distill. Beyond the $m_{0}$ values reported in Fig. 8, end-to-end entanglement generation time increases so much in the simulations that we are unable to find any reasonable estimate of an average value. Thus, for all practical purposes, the waiting time is essentially infinite. This is also indicated by the average ages approaching $m^{\star}$.
Some estimates of the $m_{0}$-thresholds up to which the two distillation policies are effective can be made using some rather simple arguments. If the fresh elementary links have age $m_{0}$, the age of the distilled link saturates to a minimum value $m^{d}_{\min}(m_{0})$ which is a function of $m_{0}$ (see, e.g., the pumping policy in Appendix C). In order to get an end-to-end link via entanglement swapping of such distilled links, the following condition must be satisfied:
$(n-1)\cdot m^{d}_{\min}(m_{0})\leq m^{\star}.$ (5)
For the configuration in Fig. 8, the above condition translates to $m^{d}_{\min}\lessapprox 4$, which in terms of $m_{0}$ is $m_{0}\leq 5$, and this is the value around which we see that the average waiting time diverges.
In the case of swap-distill, a typical trajectory to get an end-to-end link would be to first perform $(n-2)$ swaps to get end-to-end links of age $(n-1)m_{0}$, which are then distilled to links of age $m^{d}_{\min}((n-1)\cdot m_{0})$. The threshold condition thus takes the following form:
$m^{d}_{\min}((n-1)\cdot m_{0}))\leq m^{\star},$ (6)
but since distillation is ineffective when both links involved are above the cutoff age, this condition is satisfied whenever
@@
This also explains why the threshold is lower for swap-distill, since $m^{d}_{\min}(m_{0})\leq m_{0}$. Again, for the configuration in Fig. 8, this translates to $m_{0}\lessapprox 4$, which agrees well with the simulation results. Furthermore, the threshold in the case of the swap-distill policy is identical to a policy without distillation.
@@
We further elaborate on our answer to the question of whether to distill first or to swap first in Fig. 9. To do this, we plot the waiting time improvement factor for distillation, which we define to be the ratio of the average waiting time without distillation to the average waiting time with distillation. Apart from the quality of fresh links (the parameter $m_{0}$), as shown in Fig. 8, the choice between the two distillation orderings also depends on the other network parameters. As an illustration, we show that for the parameters $(n,n_{ch},m^{\star},p_{\ell})$ chosen in Fig. 9, when $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. This observation can be understood in light of the analytical results in Appendix D and Fig. 8. Indeed, we observe that for low $p_{\ell}$, links are typically older when they are ready to be swapped, and we have seen in the previous discussions in this section that distill-swap performs better in this case, because long entanglement swaps can only be possible once the ages of the participating links have been reasonably reduced. Also, distilling after entanglement swapping is highly improbable to be successful because the ages are going to only increase due to the swapping. On the other hand, when links have a high $p_{\ell}$, swap-distill performs better, because we observe that such links are typically younger when they are ready to be swapped, and they are much more conducive to successful swapping. Furthermore, in this case, the non-deterministic swaps become the rate determining step, and hence having more opportunities to perform entanglement swapping is more important. Distilling links first reduces the number of such opportunities. Furthermore, we see that for large $p_{\ell}$, distillation is either not very useful or even detrimental. Note that these trends were also qualitatively seen in Fig. 6, when CC costs were accounted for.